Find the slope of the line that contains the points (–4, –2) and (1,1).
Answer by RAY100(1637)

(-4,-2),,,(1,1)
.
m= (y2-y1)/(x2-x1) = (1-(-2))/(1-(-4)) = 3/5 ,,,,slope is positive 3/5
.
a rough graph, showing the points, connecting line, and rise / run confirms the calculation.